the geometric length of a bar magnet of pole  strength 80 A.m. and magnetic moment4Am2 ?

Grade:11

1 Answers

Arun
6 years ago
geometric length = ?
pole strength (m)= 80
M = ?
as magnetic length(2l) = 5\6 geometric length
so that 2l  =  0.15m
now  M  = m.2l = 4
4 = 80 * 2I
2I = 1/20 m = 5 cm
2I = (5/6) * geometric length
Geometric length = 5*6/5 = 6 cm
Regards
Arun (askIITians forum expert)